What is the volume in dm3 of 5.5 grams carbon dioxide (CO2) at room temperature and pressure (r.t.p)?
Given the atomic masses: C= 12 and O = 16
1
Expert's answer
2018-09-13T05:23:25-0400
At first we must calculate the number of moles for carbon dioxide.
n=m/M.
M(CO2)=12+2*16=44 g/mol
n(CO2)= 5.5/44 = 0.125 moles of CO2.
We know that 1 mole of gas at r.t.p. has a volume 24dm^3. It means that 0.125 moles of CO2 will have the volume v(CO2) = 0.125*24=3 dm^3.